### v1

# Cross-cultural Inspiration Detection and Analysis in Real and LLM-generated Social Media Data

Abstract
Inspiration is linked to various positive outcomes, such as increased creativity, productivity, and happiness. Although inspiration has great potential, there has been limited effort toward identifying content that is inspiring, as opposed to just engaging or positive. Additionally, most research has concentrated on Western data, with little attention paid to other cultures. This work is the first to study cross-cultural inspiration through machine learning methods. We aim to identify and analyze real and AI-generated cross-cultural inspiring posts. To this end, we compile and make publicly available the InspAIred dataset, which consists of 2,000 real inspiring posts, 2,000 real non-inspiring posts, and 2,000 generated inspiring posts evenly distributed across India and the UK. The real posts are sourced from Reddit, while the generated posts are created using the GPT-4 model. Using this dataset, we conduct extensive computational linguistic analyses to (1) compare inspiring content across cultures, (2) compare AI-generated inspiring posts to real inspiring posts, and (3) determine if detection models can accurately distinguish between inspiring content across cultures and data sources.
Cross-cultural Inspiration Detection and Analysis
in Real and LLM-generated Social Media Data
Oana Ignat1††thanks: Oana Ignat and Gayathri Ganesh Lakshmy contributed equally to the manuscript Gayathri Ganesh Lakshmy Rada Mihalcea1 1University of Michigan - Ann Arbor, USA 2Sri Sivasubramaniya Nadar College of Engineering - Chennai, Tamil Nadu, India {oignat, mihalcea}@umich.edu gayathri2010090@ssn.edu.in
1 Introduction
Inspiration has been a part of our world for millennia, starting with ancient Greece, where Muses were responsible for delivering divine knowledge by whispering in a poet’s ear Leavitt and Leavitt (1997), all the way to today’s creativity domain, where it is still common for artists and scientists to attribute their best ideas to a higher power, independent of their own control.
According to Thrash and Elliot (2003, 2004), inspiration is a general construct that consists of three core characteristics: evocation, transcendence, and approach motivation. Evocation refers to the process of being triggered by a stimulus, either from within (such as a creative idea that comes from the subconscious) or from outside (such as a person, object, music, or nature). Transcendence allows one to perceive something beyond their usual concerns (Milyavskaya et al., 2012). Finally, an inspired person is motivated to express, transmit, or act on their inspiration (Elliot and Thrash, 2002).
Inspiration is an area of study with promising cross-disciplinary applications in creative fields (e.g., advertisement, storytelling), education, therapy, mentorship, coaching, or social media. For instance, social network recommendation systems can mitigate potential harms by showing more positive and inspiring content to users (Ignat et al., 2021). Access to inspiring content can have a positive impact on people’s lives by offering them a fresh perspective and motivating them to take action, particularly during periods of uncertainty and concern (Oleynick et al., 2014). Moreover, inspiration facilitates progress towards goals Milyavskaya et al. (2012) and increases overall well-being Thrash et al. (2010).
Despite the compelling motivation, little research has been done on the automatic identification of content that is inspiring, as opposed to simply positive Ignat et al. (2021), and no studies have been conducted on how inspiration varies across different cultures, and whether it can be automatically generated. At the same time, the impressive generative ability of LLMs can create new opportunities for automatically generating inspiring content.
In this work, we aim to address these research gaps by focusing on three main research questions. (RQ1:) How do inspiring posts compare across cultures? (RQ2:) How do AI-generated inspiring posts compare to real inspiring posts across cultures? And (RQ3:) Can detection models effectively differentiate inspiring posts across diverse cultures and data sources?
We summarize our contributions as follows. First, we share a novel dataset comprised of cross-cultural real and generated inspiring and non-inspiring posts, for a total of 2,000 real inspiring posts, 2,000 non-inspiring posts, and 2,000 LLM-generated posts, balanced across India and the UK. Second, we make use of the dataset to conduct extensive computational linguistic analyses to compare inspiring content across cultures and data sources, i.e., India vs. the UK and AI-generated vs. human-written. Finally, we explore the effectiveness of machine learning models for detecting inspiration across diverse data sources and cultures.
2 Related Work
Automatic Inspiration Detection.
There have been a limited number of research studies conducted on inspiration, with the majority of them being carried out by the psychology and sociology communities. These studies, such as the ones by Thrash and Elliot (2003, 2004) and Elliot and Thrash (2002), have established the fundamental characteristics of inspiration. Additionally, they have developed a scale to measure the frequency with which people experience inspiration in their daily lives. These studies found that individuals who are inspired tend to be more open to new experiences and show greater absorption in their tasks. They are also more intrinsically motivated, have a strong drive to master their work, and are less competitive.
The work most similar to ours is from Ignat et al. (2021), who are the first to study inspiration through machine learning methods. To facilitate research in this domain, they release a weakly labeled dataset consisting of inspiring and non-inspiring posts collected from Reddit and annotate these posts with their effect on the reader and the emotions they transmit. They also provide a RoBERTa (Liu et al., 2019) classifier fine-tuned on human labels to provide a strong baseline for determining whether a post is inspiring. Finally, they perform extensive data analyses to gain insights into which topics inspire readers and how they influence them. Our work builds on Ignat et al. (2021) by extending it to other cultures, India and the UK, and by collecting AI-generated inspiring posts in order to compare them to human-written posts.
Human vs. LLM-generated cross-cultural text.
With the rapid development of LLMs, these models demonstrate remarkable proficiency in generating human-like text across multiple languages and styles Wu et al. (2023); Tang et al. (2023).
In particular, LLMs excel at creative writing, such as story generation (Yuan et al., 2022), advertising slogan creation (Murakami et al., 2023), and news composition (Yanagi et al., 2020). Tools like Yuan et al. (2022) can help users in their creative pursuits. By generating inspiring content, our work provides a more indirect approach, with the same final goal of helping the user express, transmit, and act on their insight.
More similar to our work, LLMs have also started to be applied to motivate people Cox et al. (2023). For example, Karinshak et al. (2023) used GPT-3 to generate messages to persuade people to receive the Covid-19 vaccine. To the best of our knowledge, we are the first to generate inspiring content using LLMs. Our work is also part of the emerging work on modeling cultural factors in LLMs (Huang and Yang, 2023; Fung et al., 2022; Ramezani and Xu, 2023). Inspiration varies across cultures. Therefore, we test the cultural knowledge of LLMs about inspiration in India and the UK, and compare it to inspiring Reddit posts from users in these countries.
Computational Linguistics for Social Media Analysis.
The advent of computational linguistics techniques has enabled researchers to analyze vast amounts of social media data for various purposes, including sentiment analysis, topic modeling, and linguistic variation across cultures (Pennebaker et al., 2007; Pang et al., 2008; Imran et al., 2020). These works have facilitated the extraction of meaningful insights from diverse linguistic contexts, paving the way for studies on cross-cultural communication in online environments.
Social media is a key source of inspiring content for younger audiences (Raney et al., 2018). Features such as hope and appreciation of beauty and excellence trigger self-transcendent emotions in videos tagged with “inspiration” on YouTube (Dale et al., 2017), as well as in #inspiring and #meaningful Tumblr memes and Facebook posts (Rieger and Klimmt, 2019; Dale et al., 2020). Similarly, we collect Reddit posts from subreddits related to inspiration across UK and Indian cultures and analyze them using computational linguistic tools such as Pennebaker et al. (2007).
3 The InspAIred Dataset
To answer our research questions, we compile a novel dataset, which we refer to as InspAIred - AI-generated Inspiring Reddit Content. Our dataset contains inspiring and non-inspiring posts from India and the UK, from two different sources: (1) crawled from Reddit and (2) generated by an LLM.
[…]
We collect 2,000 weakly labeled inspiring posts and 2,000 weakly labeled non-inspiring posts, balanced across India and the UK. We describe our data collection and annotation process below.
Data Collection.
We scrape around 5,300 posts from Reddit, a popular online platform, specifically focusing on culturally inspiring content. Following the data collection process from Ignat et al. (2021), we conduct searches within culture-related or discussion-related flairs of the subreddits using keywords, such as “inspiration” and “motivation”, to identify the relevant data in the form of both posts and comments.
More specifically, for Indian data, we primarily target the regions of Kerala, Karnataka, Maharashtra, and Tamil Nadu. Besides the general “r/india”, we also explore subreddits at the state level, such as “r/Kerala”, “r/Karnataka”, “r/Maharashtra”, and “r/TamilNadu”, which serve as hubs for focused discussions on regional culture, traditions, and social issues. Finally, we also examine subreddits from specific cities, including capital cities like Chennai and Bangalore, to capture more local perspectives and experiences. For the UK data, we follow a similar strategy to collect Reddit posts, targeting both state level subreddits, such as “r/UnitedKingdom”, as well as regional subreddits representing areas and capital cities, such as “r/London”.
We group the collected posts into two main categories: inspiring posts from India and inspiring posts from the UK. Initially, we attempted to create a more fine-grained split at the region or city level, but we faced difficulties in finding annotators from those specific regions. However, we encourage future research to explore this direction and investigate how inspiration varies within a country or within a region, along with exploring other demographic information such as language, age, gender, or income.
Data Filtering.
The collected posts are further filtered with an inspiration classifier model. We use the XLM-RoBERTa base Conneau et al. (2019) from HuggingFace.111https://huggingface.co/FacebookAI/xlm-roberta-base The model is a multilingual version of RoBERTa Liu et al. (2019), and is pre-trained on 2.5TB of filtered CommonCrawl Wenzek et al. (2020)222We choose a multilingual model, as the Indian data contains code-mix data.. We fine-tune the model for five epochs using the dataset described below, with a learning rate of and a batch size of . More implementation details can be found in the Section A.1.
For fine-tuning, we use the data from Ignat et al. (2021). The data contains around 12,000 annotated “general”, i.e., no target culture, balanced Reddit inspiring and non-inspiring posts. More details about the fine-tuning data can be found in Ignat et al. (2021). Next, we evaluate the quality of the fine-tuned model predictions by annotating a subset of the posts and comparing them with the model predictions.
Data Annotation.
The posts are annotated by crowd-sourced workers from India and the UK. For each country, we select a sample of 200/ 1,000 posts equally sampled from inspiring and non-inspiring predicted posts, to be labeled by three annotators from that country.
We use Potato Pei et al. (2022) to create the user interface and store the data. Following the guidelines from Ignat et al. (2021), the user interface contains a definition of inspiration and examples of inspiring and non-inspiring posts, as seen in Figure 2. To find and hire UK annotators, we connect Potato to Prolific.333www.prolific.com We select annotators who have the following qualifications: an approval rate >98%, living in the UK, and a right to vote (as a proxy for age). The annotators independently label 200 posts each, categorizing the posts based on their subjective judgment of whether they found the content inspiring or not. We conduct a similar annotation process with three reliable Indian annotators to label the data from India. To ensure consistency amongst the annotators, they are provided with the same interface as used for the UK annotators and shown in Figure 2.
Finally, we compute the agreement score between the annotators using the Fleiss Kappa measure Fleiss and Cohen (1973). We obtain a score of for UK data and for Indian data, indicating a fair agreement, expected for a subjective task as inspiration detection.
Given that each post is annotated by three annotators, following Ignat et al. (2021), we mark a post as inspiring if at least one annotator labeled it as inspiring.
[…]
| Annotated | Weakly-Labeled | Annotated | Weakly-Labeled | |
[…]
We find that the predictions made by the fine-tuned model are quite similar to human annotations. For Indian data, the accuracy rate is , and the F1 score is . Similarly, for UK data, the accuracy rate is , and the F1 score is . Therefore, we decide to not annotate more data and instead use the model predictions. Additionally, we further fine-tune the model on the annotated data and use it to collect more weakly labeled data. This approach aims to leverage the insights gained from the initial fine-tuning process and apply them to the user-annotated subset, thereby refining the model to better capture the nuances present in this specific dataset. We use the fine-tuned model predictions and the annotated subset as our final labeled data.
Quality Assurance.
We remove posts that are classified as toxic or hate speech using two fine-tuned XLM-RoBERTa classifiers from HuggingFace.444https://huggingface.co/textdetox/xlmr-large-toxicity-classifier555https://huggingface.co/facebook/roberta-hate-speech-dynabench-r4-target We also remove profanity from posts using a profanity classifier.666https://github.com/snguyenthanh/better_profanity We acknowledge that automatic detection of profanity, toxicity, and hate speech are active research areas that have yet to be solved (Dale et al., 2021; Vidgen et al., 2021). Finally, for each country, we manually inspect 100 random posts to ensure syntactic and semantic correctness and find the remaining toxic or hate speech posts. We find that around 99% of the posts are of good quality.
Data Statistics.
The final data statistics are shown in Table 1.
[…]
We generate 2,000 inspiring posts with GPT-4,777https://platform.openai.com/docs/guides/text-generation/chat-completions-api balanced across India and the UK. Our study can be conducted with any LLM. However, we chose GPT-4 because it is one of the largest LLMs available and has been shown to effectively emulate human texts. Achiam et al. (2023)
3.2.1 Prompt Design and Robustness
GPT-4 takes as input a list of message objects, and returns an inspiring Reddit post. We use messages, which are more interactive and dynamic compared to the classical prompt style. Specifically, we use messages with three different roles: system, user or assistant.888https://help.openai.com/en/articles/7042661-chatgpt-api-transition-guide
The prompt is first formatted with a system role, which sets the behavior of the model. This is followed by a conversation between the user and assistant, in a few-shot prompting fashion. Prior work found that LLMs function better with few-shot prompts (i.e., instructions alongside example output) rather than using zero-shot prompts (with no examples). (Brown et al., 2020)
System Prompt.
We find that we can obtain high-quality responses with additional context in our prompts. Therefore, we instruct the model to be a Reddit user from either the UK or India. To ensure that the generated data is diverse and reliable, we collect five versions of system prompts with different phrasing, with one example shown below.
Imagine you’re a person from {location} and use Reddit regularly.
User-Assistant Prompts.
We design two rounds of conversations between a user and an assistant, where the user asks for a Reddit post or comment, and the assistant responds to the request. We use few-shot prompting by providing the assistant in the first round of conversation with an inspiring post, which is randomly extracted from the real posts annotated as inspiring by the majority of annotators. Finally, the answer to the second user is automatically generated by the assistant and used to collect the GPT-4 generated inspiring post. A user message is shown below.
Write a Reddit post or comment of maximum 100 tokens about what inspires you.
Quality Assurance.
To ensure the quality of our generated data, we conduct sanity checks to review approximately 200 inspiring posts balanced across the UK and India. The posts are checked for cultural knowledge, factuality, semantic and syntax errors, and style. Based on the feedback, we find that the posts are semantically and syntactically accurate, possess cultural knowledge, and are often more complex than real posts.
[…]
| ✗ What is a common meal there? I’ve always thought Indian palates to be quite spicy and rich. | ✗ N/A | |
[…]
| ✗ Mate all he does is play football and misses pens. Not that special | ✗ N/A |
4 Cross-Cultural Inspiration Analysis across Real and LLM-generated posts
In line with previous work (Jakesch et al., 2023), we also find during manual inspection that it is challenging to differentiate between LLM-generated texts and those written by humans (see Table 2). Therefore, we perform computational linguistic analyses to compare real and LLM-generated text across cultures. This section addresses our first two research questions: RQ1: How do inspiring posts compare across cultures? and RQ2: How do AI-generated inspiring posts compare to real inspiring posts across cultures?.
4.1 Stylistic and Structural Features
We assess the linguistic style and structure of real and LLM-generated inspiring posts across India and the UK in terms of (1) analytic writing, (2) descriptiveness, and (3) readability.
Analytic writing
index measures the complexity and sophistication of the writing, which can be an indicator of advanced thinking. The formula for analytic writing is from LIWC scores (Jordan et al., 2019; Pennebaker et al., 2014) More information about LIWC can be found in Section 4.2. We display the scores in Table 3. The low complexity scores of Reddit posts are primarily negative due to their high usage of pronouns and lack of articles. We find that LLM-generated inspiring posts from the UK and India are more complex than real posts, which aligns with our initial observations from data quality checks. At the same time, there is no significant difference in text complexity between real inspiring posts from India and those from the UK.
Descriptiveness
can be measured by the frequency of adjectives used in language patterns. Texts with high rates of adjectives tend to be more elaborate and narrative-like compared to texts with low rates of adjectives. (Chung and Pennebaker, 2008) In Table 3, we find that LLM-generated inspiring posts from the UK and India are more descriptive than real posts. Additionally, real inspiring posts from India are more descriptive than those from the UK.
Readability
considers not only word count, but also word complexity. For instance, longer words are more complex than shorter ones. We use the Flesch Reading Ease metric (Flesch, 1948), which counts the number of words per sentence and syllables per word. In Table 3, we find that LLM-generated inspiring posts from the UK and India are less readable than real posts. Additionally, real inspiring posts from India are less readable than those from the UK. Furthermore, when measuring post length, LLM-generated UK posts are longer than real posts and UK posts are shorter than Indian posts.
[…]
| Analytic | -16.7 18.0 | -6.9 10.0 | -17.1 19.9 | -7.0 11.4 |
| Descriptive | 8.2 7.8 | 8.9 3.6 | 6.6 5.9 | 9.1 3.9 |
| Readable | 36.3 63.1 | 12.1 20.6 | 53.2 52.2 | 29.6 18.0 |
| Word Count | 61.1 59.8 | 66.9 20.3 | 43.9 49.3 | 49.9 16.8 |
4.2 Semantic and Psycholinguistic Features
We assess the semantic and psycholinguistic differences between LLM-generated and real inspiring posts across India and the UK, using topic modeling and LIWC psycholinguistic markers.
Data Pre-processing.
We use spaCy library999https://spacy.io/models to pre-process the data: tokenize each post, lowercase tokens, remove stop words, remove numbers, symbols, emojis, links, and lemmatize tokens.
Topic Modeling.
We use Scattertext Kessler (2017), a tool used to create interactive visualizations of linguistic patterns. After initial pre-processing, we structure the text data for a better representation of information. Next, we use the n-gram representation of the text data to conduct topic extraction through sentence-level clustering. We utilize methods like Non-negative Matrix Factorization (NMF) to decompose the TF-IDF matrix to identify the latent thematic structures within the corpus Ramos et al. (2003).
Results. We analyze the n-gram and topic distributions by various dimensions: inspiring vs. non-inspiring, Indian vs. UK, and real vs. generated. We display the topic distribution across real and generated UK inspiring posts in Figure 3. Further topic and n-gram distributions are displayed in Section A.2.
vs. ✗: Comparing inspiring to non-inspiring real posts, we find that, in the Indian data (Fig. 5, 7), the topic of people is frequently discussed in an inspiring context. Moreover, the most commonly occurring theme in non-inspiring posts is bot, encompassing textual content regarding Reddit rules and moderators.
In the UK data (Fig. 6, 8), the topic of life is amongst the most commonly occurring themes in inspiring posts, with discussions surrounding career, luck, and pension. On the other hand, the non-inspiring posts contain themes like dark, including conversations about rain and winter.
[…]
In the LLM-generated UK data (Fig. 3, top left), the topic nhs is emphasized, where common words include pandemic, heroes, and staff. Additionally, discussions often revolve around adversity, with mentions of resilience, determination, and spirit. In contrast, in the real data from the UK (Fig. 3, bottom right), discussions related to job dominate, with mentions of salary, savings, and time. Moreover, there is also a significant focus on exercise, featuring words like discipline, and motivate.
LIWC psycholinguistic markers.
We use Linguistic Inquiry and Word Count (LIWC), a gold-standard text analysis tool (Pennebaker et al., 2007, 2015), to obtain the words related to human cognitive processes from each post. Specifically, we use the LIWC2015 dictionary, which contains 6,400 words and word stems, each related to a cognitive category. As an example, the word “mother” is assigned the following cognitive categories: female, family, social.
Results. We display the top 24 categories and their LIWC scores, with the most significant differences across dimensions in Table 4. Across all data, we find that male words are more frequent than female, the most common pronoun is I, the most frequent tense is present, and the most common emotion is positive.
vs. ✗: Comparing inspiring to non-inspiring real posts, we find that, in the Indian data, inspiring posts are more likely to be related to socializing, insight, feelings, perception, affection, and contain more positive emotions. Additionally, Indian inspiring posts tend to have more comparisons and use more words related to achievement, health, reward, and work.
UK real inspiring posts contain more words related to affection, comparisons, feelings, achievement, health, home, leisure, money, reward, and work than non-inspiring posts. Furthermore, compared to Indian inspiring posts, UK inspiring posts have fewer words related to family, socializing, affection, perception, religion, less positive emotions and more words related to achievement, health, home, leisure, money, rewards, and work.
[…]
UK real inspiring posts contain more words related to comparisons, feelings, health, home, leisure, money, rewards and work than LLM-generated UK posts. Conversely, real UK posts contain fewer words related to socializing, affection, insight, perception, achievement and religion than LLM-generated UK posts. Furthermore, compared to LLM-generated Indian posts, LLM-generated UK posts have fewer words related to socializing, achievement, leisure, money, religion, work and more words related to affection, perception and more positive emotions.
[…]
| Social Processes | ||||||
| FAMILY | 0.4 | 0.4 | 0.2 | 0.2 | 0.1 | 0.2 |
| FRIEND | 0.3 | 0.4 | 0.2 | 0.2 | 0.4 | 0.2 |
| FEMALE | 0.4 | 0.3 | 0.2 | 0.4 | 0.4 | 0.2 |
| MALE | 1.1 | 1.0 | 0.9 | 0.9 | 1.3 | 1.7 |
| SOCIAL | 9.3 | 7.7 | 8.6 | 7.4 | 8.1 | 8.1 |
| Affective Processes | ||||||
| AFFECT | 6.1 | 4.3 | 8.4 | 5.4 | 4.8 | 9.4 |
| NEGEMO | 1.4 | 1.3 | 1.0 | 1.3 | 1.6 | 0.9 |
| POSEMO | 4.5 | 3.0 | 7.3 | 4.1 | 3.1 | 8.3 |
| Cognitive Processes | ||||||
| COMPARE | 3.0 | 2.4 | 1.8 | 3.0 | 2.7 | 1.8 |
| INSIGHT | 2.3 | 2.0 | 4.8 | 2.2 | 2.1 | 4.9 |
| Perceptual Processes | ||||||
| FEEL | 0.6 | 0.4 | 0.4 | 0.7 | 0.3 | 0.4 |
| PERCEPT | 2.4 | 2.0 | 2.1 | 2.1 | 2.1 | 2.4 |
| Personal Concerns | ||||||
| ACHIEV | 1.9 | 1.0 | 3.5 | 2.4 | 1.6 | 2.6 |
| HEALTH | 0.7 | 0.4 | 0.9 | 1.0 | 0.8 | 0.8 |
| HOME | 0.3 | 0.2 | 0.1 | 0.7 | 0.4 | 0.2 |
| LEISURE | 1.5 | 1.4 | 1.2 | 1.7 | 1.4 | 0.9 |
| MONEY | 0.7 | 0.8 | 0.6 | 1.2 | 0.8 | 0.4 |
| RELIG | 0.4 | 0.5 | 1.1 | 0.1 | 0.1 | 0.6 |
| REWARD | 2.0 | 1.1 | 1.0 | 2.2 | 1.6 | 1.1 |
| WORK | 3.0 | 1.9 | 3.0 | 3.9 | 2.7 | 1.9 |
5 Cross-Cultural Inspiration Detection across Real and LLM-generated posts
To answer our last research question – RQ3: Can detection models effectively differentiate inspiring posts across diverse cultures and data sources? – we fine-tune a multi-label classification model to identify if a post is inspiring, represents India or the UK, and whether it is real or LLM-generated.
Implementation Details.
We use the XLM-RoBERTa base model Conneau et al. (2019), pre-trained on 2.5TB of filtered CommonCrawl Wenzek et al. (2020). We fine-tune the model on our dataset for five epochs with a learning rate of and a batch size of , while monitoring the validation performance and selecting the best model checkpoint based on the F1 metric.
Model Training Setup.
We experiment with two setups to split the data into training, validation, and test sets: a default train-val-test split of 64-16-20% and a few-shot split of 8-2-90%.
Results.
Since the dataset is evenly distributed across classes, a random baseline results in a 50% accuracy score.
The XLM-RoBERTa model with few-shot setup achieves an accuracy of and F1 score of across all labels, while the default setup achieves an accuracy of and F1 score of . Figure 4 displays the disaggregated results for each label. We find that even with very few training data (600 posts), in the few-shot setup, the XLM-RoBERTa model learns to accurately distinguish inspiring content across cultures (India, UK) and data sources (real and generated). Furthermore, we do not find significant differences in performance across cultures and data sources. Finally, the model can accurately distinguish between real and generated inspiring posts in both few-shot and default training setups.
6 Conclusion
In this paper, we introduced the task of cross-cultural inspiration detection and generation in social media data. To facilitate research in this domain, we released InspAIred, a dataset of 2,000 real inspiring posts, 2,000 real non-inspiring posts, and 2,000 LLM-generated inspiring posts, evenly distributed across India and the UK. We performed extensive linguistic data analyses to gain insight into what topics inspire each culture and compare AI-generated inspiring posts to real inspiring posts across various linguistic dimensions. Despite the difficulty humans have in distinguishing between real posts and those generated by LLMs, we found that these posts have noticeable differences in style, structure, and semantics and that, even with little data, fine-tuned models accurately distinguish inspiring content across cultures and data sources.
We hope our work will enable the exploration of various applications to improve creativity and motivation, including storytelling, advertising, and social media, as well as therapy and coaching. Our dataset can be used to test, fine-tune, and analyze other models, and it is publicly available at https://github.com/MichiganNLP/cross_inspiration.
Limitations
1. A more fine-grained data split
Our posts are divided into two main categories - inspiring posts from India and inspiring posts from the UK. Initially, we tried to create a more detailed classification at the region or city level, but we faced difficulties in finding annotators from those specific regions. However, we encourage future research to explore this direction and investigate how inspiration varies within a country or region, as well as explore other demographic information, such as language, age, gender, or income.
Study case for India.
The extensive discussions on Indian subreddits, like r/AskIndia and r/India, offer straightforward access to text-based motivational narratives. Additionally, region-based subReddits, like r/Chennai and r/Mumbai, provide diverse insights into localized experiences and discussions. The nature of posts on this topic is mostly anecdotal. Reddit users contribute by sharing personal experiences that have inspired them.
2. Limited number of cross-cultural inspiring posts
It is crucial to recognize that inspiration is not always explicitly articulated through words like “inspiring” or “motivation”, especially within cultures that aren’t as open about talking about such topics. In many instances, it manifests more subtly, embedded within narratives, imagery, or cultural expressions. This implicit nature of inspiration adds another layer of complexity to the data collection process.
The language barrier also introduces an additional barrier, as it requires linguistic fluency and cultural understanding to interpret and analyze inspiring content effectively. Posts in languages other than English may contain nuances and cultural references that are not easily translatable. Moreover, it becomes more challenging to collect inspiring content when data collection is restricted to certain countries. That is why we collected fewer posts compared to Ignat et al.’s “general” inspiring posts.
3. Relevance of LLM-based data to current times
In our reliance on LLM-based data, it’s imperative to recognize the temporal limitations inherent in its training corpus. While these models offer remarkable capabilities in understanding and generating text, they might not fully capture the current cultural and societal context in their outputs. This limitation can be primarily attributed to the fact that these models are trained on data only up to a specific cutoff date, i.e., GPT-4 only learned from data dated up to September 2021.
Appendix A Appendix
A.1 Fine-tuning Process
XLM-RoBERTa Model.
Implementation Details. We experiment with three different fine-tuned classification models for the purpose of weakly labeling the presence of inspiration in the dataset. The first attempt involved fine-tuning the model on the general inspiration dataset introduced in Ignat et al. (2021). In each of the experiments, the data is split into three subsets: training, validation, and test sets, using an 80:10:10 ratio. The fine-tuning process involved training the model for 5 epochs with a learning rate of and a batch size of . We monitored the model performance on a separate validation set and selected the best model checkpoint based on accuracy metric.
In the second experiment, we focused on a subset of the dataset used in annotation, consisting of 200 posts. These posts were labeled as inspiring if at least one user considered them to be so. We fine-tuned the base XLM-RoBERTa model using this subset, maintaining the same training configurations as in the initial experiment. In the last experiment, we leveraged the user-annotated posts to further refine the model trained on the general inspiration dataset from the first experiment. This approach aimed to leverage the insights gained from the initial fine-tuning process and apply them to the user-annotated subset, thereby refining the model to better capture the nuances present in this specific dataset.

### v2

@@
 We aim to identify and analyze real and AI-generated cross-cultural inspiring posts.
-To this end, we compile and make publicly available the InspAIred dataset, which consists of 2,000 real inspiring posts, 2,000 real non-inspiring posts, and 2,000 generated inspiring posts evenly distributed across India and the UK.
+To this end, we make publicly available the InspAIred dataset, which consists of 2,000 real inspiring posts, 2,000 real non-inspiring posts, and 2,000 generated inspiring posts evenly distributed across India and the UK.
 The real posts are sourced from Reddit, while the generated posts are created using the GPT-4 model.
 Using this dataset, we conduct extensive computational linguistic analyses to (1) compare inspiring content across cultures, (2) compare AI-generated inspiring posts to real inspiring posts, and (3) determine if detection models can accurately distinguish between inspiring content across cultures and data sources.
+Our dataset can be accessed alongside our classification models at https://github.com/MichiganNLP/cross_inspiration.
 Cross-cultural Inspiration Detection and Analysis
@@
 Despite the compelling motivation, little research has been done on the automatic identification of content that is inspiring, as opposed to simply positive Ignat et al.
-(2021), and no studies have been conducted on how inspiration varies across different cultures, and whether it can be automatically generated.
-At the same time, the impressive generative ability of LLMs can create new opportunities for automatically generating inspiring content.
+(2021), and no studies have been conducted on how inspiration varies across cultures, and whether it can be automatically generated.
+At the same time, LLMs’ impressive generative ability can create new opportunities for automatically generating inspiring content.
 In this work, we aim to address these research gaps by focusing on three main research questions.
@@
 (RQ2:) How do AI-generated inspiring posts compare to real inspiring posts across cultures?
-And (RQ3:) Can detection models effectively differentiate inspiring posts across diverse cultures and data sources?
+and (RQ3:) Can detection models differentiate inspiring posts across cultures and data sources?
 We summarize our contributions as follows.
 First, we share a novel dataset comprised of cross-cultural real and generated inspiring and non-inspiring posts, for a total of 2,000 real inspiring posts, 2,000 non-inspiring posts, and 2,000 LLM-generated posts, balanced across India and the UK.
-Second, we make use of the dataset to conduct extensive computational linguistic analyses to compare inspiring content across cultures and data sources, i.e., India vs.
+Second, we make use of the dataset to conduct computational linguistic analyses to compare inspiring content across cultures and data sources, i.e., India vs.
 the UK and AI-generated vs.
@@
 These studies, such as the ones by Thrash and Elliot (2003, 2004) and Elliot and Thrash (2002), have established the fundamental characteristics of inspiration.
-Additionally, they have developed a scale to measure the frequency with which people experience inspiration in their daily lives.
+Additionally, they have developed a scale to measure the frequency with which people experience inspiration in their lives.
 These studies found that individuals who are inspired tend to be more open to new experiences and show greater absorption in their tasks.
@@
 They also provide a RoBERTa (Liu et al., 2019) classifier fine-tuned on human labels to provide a strong baseline for determining whether a post is inspiring.
-Finally, they perform extensive data analyses to gain insights into which topics inspire readers and how they influence them.
+Finally, they perform data analyses to gain insights into which topics inspire readers and how they influence them.
 Our work builds on Ignat et al.
@@
 (2022) can help users in their creative pursuits.
-By generating inspiring content, our work provides a more indirect approach, with the same final goal of helping the user express, transmit, and act on their insight.
+By generating inspiring content, our work is more indirect but still aims to help the user express and act on their insights.
 More similar to our work, LLMs have also started to be applied to motivate people Cox et al.
@@
 (2021), the user interface contains a definition of inspiration and examples of inspiring and non-inspiring posts, as seen in Figure 2.
-To find and hire UK annotators, we connect Potato to Prolific.333www.prolific.com We select annotators who have the following qualifications: an approval rate >98%, living in the UK, and a right to vote (as a proxy for age).
+To find and hire UK annotators, we connect Potato to Prolific.333www.prolific.com We select three annotators who have the following qualifications: an approval rate >98%, living in the UK, and a right to vote (as a proxy for age).
 The annotators independently label 200 posts each, categorizing the posts based on their subjective judgment of whether they found the content inspiring or not.
@@
 Finally, we compute the agreement score between the annotators using the Fleiss Kappa measure Fleiss and Cohen (1973).
-We obtain a score of for UK data and for Indian data, indicating a fair agreement, expected for a subjective task as inspiration detection.
-Given that each post is annotated by three annotators, following Ignat et al. (2021), we mark a post as inspiring if at least one annotator labeled it as inspiring.
-[…]
-| Annotated | Weakly-Labeled | Annotated | Weakly-Labeled | |
-[…]
-[…]
+We obtain a score of for UK data and for Indian data, indicating a fair agreement, aligning with Ignat et al.
+(2021) who obtained a agreement score.
+Furthermore, by manually examining the low-agreement examples, we find that inspiration is highly subjective and personal, relying on one’s beliefs, interests, or upbringing (e.g., someone not interested in nature will probably not be inspired by David Attenborough).
+Finally, we make available all the annotations to encourage future work to research the general or personal nature of inspiration based on data and downstream applications.
+Given that each post is annotated by three annotators, following Ignat et al.
+(2021), we mark a post as inspiring if at least one annotator labeled it as inspiring.
+We do not select the majority label to avoid excluding other annotators’ opinions.
+Disagreement among annotators in subjective tasks may reflect systematic differences in opinions across groups, not noise (Fleisig et al., 2023).
 We find that the predictions made by the fine-tuned model are quite similar to human annotations.
@@
 The final data statistics are shown in Table 1.
+[…]
+| Annotated | Weakly-Labeled | Annotated | Weakly-Labeled | |
+[…]
+[…]
 […]
@@
 Based on the feedback, we find that the posts are semantically and syntactically accurate, possess cultural knowledge, and are often more complex than real posts.
+More information is provided in the Section A.1.
 […]
@@
 | ✗ Mate all he does is play football and misses pens. Not that special | ✗ N/A |
+Cost.
+We generated 2,000 posts for two cultures, for a total cost of around 50$. (0.03 per 1K input tokens and 0.06 per 1K output tokens).
 4 Cross-Cultural Inspiration Analysis across Real and LLM-generated Posts
@@
 We use the Flesch Reading Ease metric (Flesch, 1948), which counts the number of words per sentence and syllables per word.
-In Table 3, we find that LLM-generated inspiring posts from the UK and India are less readable than real posts.
+In Table 3, we find that LLM-generated inspiring posts are less readable than real posts.
 Additionally, real inspiring posts from India are less readable than those from the UK.
@@
 Topic Modeling.
-We use Scattertext Kessler (2017), a tool used to create interactive visualizations of linguistic patterns.
-After initial pre-processing, we structure the text data for a better representation of information.
-Next, we use the n-gram representation of the text data to conduct topic extraction through sentence-level clustering.
-We utilize methods like Non-negative Matrix Factorization (NMF) to decompose the TF-IDF matrix to identify the latent thematic structures within the corpus Ramos et al.
-(2003).
+We use Scattertext Kessler (2017) to create interactive visualizations of linguistic patterns.
+We use the n-gram representation of the text data to conduct topic extraction through sentence-level clustering.
+Ramos et al.
+(2003)
 Results.
@@
 We display the topic distribution across real and generated UK inspiring posts in Figure 3.
-Further topic and n-gram distributions are displayed in Section A.2.
+Further topic and n-gram distributions are in Section A.3.
 vs.
@@
 5, 7), the topic of people is frequently discussed in an inspiring context.
-Moreover, the most commonly occurring theme in non-inspiring posts is bot, encompassing textual content regarding Reddit rules and moderators.
+Moreover, the most commonly occurring theme in non-inspiring posts is bot, encompassing content regarding Reddit rules.
 In the UK data (Fig.
@@
 LIWC Psycholinguistic Markers.
-We use Linguistic Inquiry and Word Count (LIWC), a gold-standard text analysis tool (Pennebaker et al., 2007, 2015), to obtain the words related to human cognitive processes from each post.
+We use Linguistic Inquiry and Word Count (LIWC), a popular text analysis tool (Pennebaker et al., 2007, 2015), to obtain the words related to human cognitive processes from each post.
 Specifically, we use the LIWC2015 dictionary, which contains 6,400 words and word stems, each related to a cognitive category.
 As an example, the word “mother” is assigned the following cognitive categories: female, family, social.
+Even though it is less biased than embedding-based models, LIWC has limitations which we discuss in the Limitation section.
 Results.
 We display the top 24 categories and their LIWC scores, with the most significant differences across dimensions in Table 4.
-Across all data, we find that male words are more frequent than female, the most common pronoun is I, the most frequent tense is present, and the most common emotion is positive.
+We find that male words are more frequent than female, the most common pronoun is I, the most frequent tense is present, and the most common emotion is positive.
 vs.
-✗: Comparing inspiring to non-inspiring real posts, we find that, in the Indian data, inspiring posts are more likely to be related to socializing, insight, feelings, perception, affection, and contain more positive emotions.
+✗: Comparing inspiring to non-inspiring real posts, we find that the Indian inspiring posts are more related to socializing, insight, feelings, perception, affection, and contain more positive emotions.
 Additionally, Indian inspiring posts tend to have more comparisons and use more words related to achievement, health, reward, and work.
@@
 To answer our last research question – RQ3: Can detection models effectively differentiate inspiring posts across diverse cultures and data sources?
-– we fine-tune a multi-label classification model to identify if a post is inspiring, represents India or the UK, and whether it is real or LLM-generated.
+– we fine-tune several classification models to identify if a post is inspiring, represents India or the UK, and whether it is real or LLM-generated.
 Implementation Details.
-We use the XLM-RoBERTa base model Conneau et al.
+As a baseline, we use a Random Forest classifier (Ho, 1995) with the default settings from sklearn (Pedregosa et al., 2011) and TF-IDF Ramos et al. (2003) as features.
+As a main model, we use the XLM-RoBERTa base model Conneau et al.
 (2019), pre-trained on 2.5TB of filtered CommonCrawl Wenzek et al.
 (2020).
-We fine-tune the model on our dataset for five epochs with a learning rate of and a batch size of , while monitoring the validation performance and selecting the best model checkpoint based on the F1 metric.
+We fine-tune the model on our dataset for five epochs with a learning rate of and a batch size of .
+Finally, we also use the pre-trained Llama 2.7b model, from HuggingFace.101010https://huggingface.co/meta-llama/Llama-2-7b We fine-tune the model on our dataset using LoRA Hu et al.
+(2022) for five epochs with a rank of , learning rate of and a batch size of .111111All the model parameters are shared in the repository.
 Model Training Setup.
@@
 Since the dataset is evenly distributed across classes, a random baseline results in a 50% accuracy score.
-The XLM-RoBERTa model with few-shot setup achieves an accuracy of and F1 score of across all labels, while the default setup achieves an accuracy of and F1 score of .
 Figure 4 displays the disaggregated results for each label.
-We find that even with very few training data (600 posts), in the few-shot setup, the XLM-RoBERTa model learns to accurately distinguish inspiring content across cultures (India, UK) and data sources (real and generated).
-Furthermore, we do not find significant differences in performance across cultures and data sources.
-Finally, the model can accurately distinguish between real and generated inspiring posts in both few-shot and default training setups.
+The Random Forest baseline with TF-IDF scores, trained in a default setup achieves high accuracy scores (), suggesting that important words and topics contribute significantly to inspiration detection across culture and source.
+This finding is also supported by the topic modeling results in Section 4.2, which shows different topic distributions for each class.
+Given that Random Forest is an interpretable model, we encourage future work to further analyse the importance of different features for inspiration detection.
+Finally, we find that even with very few training data (600 posts), in the few-shot setup, the main models, Llama 2.7b and XLM-RoBERTa, learn to accurately distinguish inspiring content across cultures (India, UK) and data sources (real and generated), without significant differences in performance across cultures and data sources.
 6 Conclusion
@@
 Despite the difficulty humans have in distinguishing between real posts and those generated by LLMs, we found that these posts have noticeable differences in style, structure, and semantics and that, even with little data, fine-tuned models accurately distinguish inspiring content across cultures and data sources.
-We hope our work will enable the exploration of various applications to improve creativity and motivation, including storytelling, advertising, and social media, as well as therapy and coaching.
-Our dataset can be used to test, fine-tune, and analyze other models, and it is publicly available at https://github.com/MichiganNLP/cross_inspiration.
+We hope our work will enable the exploration of various applications to improve creativity and motivation, including storytelling, advertising, social media, therapy, and coaching.
+Our dataset is available to test, fine-tune, and analyze other models, and it can be accessed alongside our classification models athttps://github.com/MichiganNLP/cross_inspiration.
 Limitations
-1. A more fine-grained data split
+A more fine-grained data split.
 Our posts are divided into two main categories - inspiring posts from India and inspiring posts from the UK.
@@
 Reddit users contribute by sharing personal experiences that have inspired them.
-2. Limited number of cross-cultural inspiring posts
+A limited number of cross-cultural inspiring posts.
 It is crucial to recognize that inspiration is not always explicitly articulated through words like “inspiring” or “motivation”, especially within cultures that aren’t as open about talking about such topics.
@@
 That is why we collected fewer posts compared to Ignat et al.’s “general” inspiring posts.
-3. Relevance of LLM-based data to current times
+Data generated with a closed-source model.
+We use GPT-4 to generate the inspiring posts, which is not an open-source LLM.
+We believe it would be valuable for future work to consider generating more data using open-source models like LLaMA or BLOOM.
+At the same time, we publicly released the data generated with GPT-4 so that others can build on this dataset, and chose this model due to its SOTA performance, worldwide accessibility, and popularity.
+Relevance of LLM-based data to current times.
 In our reliance on LLM-based data, it’s imperative to recognize the temporal limitations inherent in its training corpus.
 While these models offer remarkable capabilities in understanding and generating text, they might not fully capture the current cultural and societal context in their outputs.
-This limitation can be primarily attributed to the fact that these models are trained on data only up to a specific cutoff date, i.e., GPT-4 only learned from data dated up to September 2021.
+This limitation can be primarily attributed to the fact that these models are trained on data only up to a specific cutoff date, i.e., GPT-4 only learned from data dated up to January 2022.
+Limitations of the LIWC-based data analysis.
+LIWC is a dictionary where each word is associated with a psycholinguistics category.
+Therefore, it is less biased than embedding-based models that often contain biased word associations (e.g., man-programmer, woman-homemaker).
+Based on a simple and scalable word count approach, LIWC is one of the most commonly used data analysis tools.
+However, it has several limitations.
+First, it does not consider word sense disambiguation based on context, which can lead to inaccurate results Schwartz et al.
+(2013).
+Second, the model was developed and validated for the analysis of long-form writing, which may impact its validity on short social media posts Panger (2016).
+Third, ongoing research is being conducted to measure the extent to which LIWC demonstrates cross-linguistic and cross-cultural generalizability, given that most studies have been conducted in English Dudău and Sava (2021).
 Appendix A Appendix
-A.1 Fine-tuning Process
+A.1 Data Quality Assurance
+We mention in the paper that the posts are checked for cultural knowledge, factuality, semantic and syntax errors, and style.
+Specifically, the authors each check 200 real and generated posts from each culture.
+One author is from India and is, therefore, able to authenticate the cultural and factual aspects of the sampled posts (e.g., Sal Khan and his education contributions, the ISRO, and its scientific progress in India) and check the correctness of the few (<10% of the data) code switch posts between English and Tamil.
+All the authors possess cultural knowledge about the UK, to recognize the cultural and factual elements of the posts.
+When unsure, they manually searched Wikipedia for the factuality of the posts (e.g., Sir David Attenborough’s contributions to nature, the British Nicola Adams, the first-ever female boxer to retain an Olympic title).
+Finally, all the authors possess proficient English skills to check for semantic and syntax errors in the sampled posts.
+We find that apart from a few typos (only in real posts), very common in social media posts, the vast majority of the posts (98%) are correct.
+In terms of style, we observe that LLM-generated sampled posts are more formal and complex than real posts (e.g., there are no typos, and there are more articles and prepositions).
+Our analysis also confirmed this by automatically measuring the analytic index.
+A.2 Fine-tuning Process
 XLM-RoBERTa Model.
@@
 This approach aimed to leverage the insights gained from the initial fine-tuning process and apply them to the user-annotated subset, thereby refining the model to better capture the nuances present in this specific dataset.
+A.3 Topic Analysis